# Meta-Homogenization for Knitwear Simulation

### v2

@@ Abstract.
This paper presents meta-homogenization, a spatially varying homogenization scheme for knitwear simulation. We are motivated by the observation that macro-scale fabric dynamics is strongly correlated with its underlying knitting patterns. Therefore, homogenization towards a single material is less effective when the knitting is complex and non-repetitive. Our method tackles this challenge by homogenizing the yarn-level material locally at volumetric elements. Assigning a virtual volume of a knitting structure enables us to model bending and twisting effects via a simple volume-preserving penalty and thus effectively alleviates the material nonlinearity. We employ an adjoint Gauss-Newton formulation to battle the dimensionality challenge of such per-element material optimization. This intuitive material model makes the forward simulation GPU-friendly. To this end, our pipeline also equips a novel domain-decomposed subspace solver crafted for GPU projective dynamics, which makes our simulator hundreds of times faster than the yarn-level simulator. Experiments validate the capability and effectiveness of meta-homogenization. Our method produces realistic animations of knitwear matching the quality of full-scale yarn-level simulations. It is also orders of magnitude faster than existing homogenization techniques in both the training and simulation stages.
@@
Figure 1. Yangge dance.  We propose a meta-homogenization algorithm for knitwear simulation. Our pipeline takes full-scale yarn-level simulation observations as input and learns a homogenization out of the input poses. We name this procedure meta-homogenization because we do not homogenize yarn-level dynamics to a codimensional sheet/shell domain but to a volumetric enclosure. Doing so allows us to use an intuitive hyperelastic material model with a volume-preserving constraint to capture nonlinear behaviors like bending and twisting. Meta-homogenization varies spatially. Each volumetric element has its own material parameter, which substantially enhances the expressivity of our model. This is a challenging problem, and our pipeline includes several computation-wise contributions such as adjoint Gauss-Newton, sample-based optimization, and a novel domain-decomposed projective dynamics solver. In the teaser figure, we show snapshots of an animated virtual character performing a Yangge dance, where we simulate a homogenized tetrahedron mesh with about $390$K elements that encapsulates a sweater. Cable patterns on the garment contribute to an irregular and sophisticated material distribution (as visualized in the sub-figure). In this example, the mesh-level simulation runs at $604$ ms per frame on average under $\Delta t=1/150$ sec. The yarn-level sweater model, on the other hand, consists of over $3$M DOFs. Our meta-homogenization makes the simulation 100$\sim$ 100$\times$ 100 × faster than simulating the sweater at the yarn level.
@@ 1. Introduction
We propose an entirely different perspective for efficient simulation of knits that can even handle non-repeating knitting patterns. In contrast to traditional material homogenization methods or HYLC, we introduce a novel approach: meta-homogenization. We focus on studying the heterogeneity, i.e., the spatial variation of material parameters while keeping the constitutional relation simple. In particular, we incorporate a virtual volume representing the yarn material and exploit spatially varying volume-preserving penalties to capture desired knit characteristics, such as stretching, shearing, and bending, without over-complicating the material model.
We discretize the fabric at the yarn level and embed it within a volumetric mesh. This mesh is of high resolution to capture the local structural characteristics of yarn loops. We learn a macro-scale material model over this volumetric space, aligning it with dynamic YLS results. In that respect, our method can also be understood as applying local homogenization at each mesh element. Our rationale resembles curve fitting, where using multiple low-degree polynomials (e.g., splines) is often preferred over employing a single high-degree polynomial. Likewise, a composite material model, despite its simplicity, offers a more versatile design space compared to a homogenized material, since each volumetric element possesses independent material freedoms. We introduce an adjoint Gauss-Newton method to regress the yarn material via a high-dimension space-time optimization problem. The simplicity of the material model makes the meta homogenization well-suited for efficient forward GPU solvers such as projective dynamics (PD) (Bouaziz et al., 2014). To this end, we devise a domain-decomposed multi-level solver for the global stage of PD specially crafted for our runtime pipeline.
@@ 2. Related work
Modern fabrics and garments are fabricated through sophisticated workmanship and exhibit intricate material response. Nevertheless, modeling cloth with a sheet-like shell remains prevalent (Choi and Ko, 2005; Grinspun et al., 2003) due to its simplicity and intuitiveness. Mass-spring and/or triangle meshes offer a natural discretization of sheet-based cloth, and its equation of motion can be derived via elastic energies that measure the cloth deformation (Terzopoulos et al., 1987). Implicit integration has become a standard component since the seminal work of Baraff and Witkin (1998). Unlike a rubber membrane, cloth fabrics are less stretchable but tend to be bent more easily. This feature inspires various techniques to enforce the inextensibility (Provot, 1995; Goldenthal et al., 2007; English and Bridson, 2008; Wang et al., 2010) with numerical stability. Cloth bending, on the other hand, can be parameterized by the dihedral angle or bending modes (Bridson et al., 2005). It is also possible to exploit isometric mesh deformations to use the so-called quadratic bending (Bergou et al., 2006; Garg et al., 2007), which possesses a constant Hessian. Kim (2020) reveals the connection between model of Baraff and Witkin (1998) and the finite element method (FEM) (Bathe, 2006). To capture detailed wrinkles and folds on the cloth, adaptive remeshing (Narain et al., 2013, 2012) or mixed discretization (Guo et al., 2018; Weidner et al., 2018) are proven to be effective. Another important aspect of sheet-level cloth animation is to handle collisions, especially self-collisions (Volino and Thalmann, 2000; Bridson et al., 2002; Baraff et al., 2003). Penalty-based methods (Bouaziz et al., 2014; Wu et al., 2020) are straightforward to implement but often with stability issues. Provot (1997); Harmon et al. (2008) introduce approaches that group multiple collisions into “impact zones” and treat them as rigid bodies, allowing for some sliding motion. Constraint-based collision handling methods represent contact as constraints based on exact Coulomb friction (Otaduy et al., 2009; Li et al., 2018).
@@
Data-driven or learning-based cloth simulation is considered an effective approach to enhance the realism of continuum shell models. Wang et al. (2011) estimate planar and bending stiffness through separate tests, while automatic devices have been developed for acquiring fabric stiffness (Miguel et al., 2012) and friction parameters (Miguel et al., 2013). Similar tensile tests have also been utilized by Clyde et al. (2017) to estimate the planar stiffness of woven fabrics. Feng et al. (2022) use a simulation-in-the-loop framework to estimate the fabric bending stiffness. ClothCap aims to reconstruct multiple garments from the 3D scans (Pons-Moll et al., 2017). Differentiable simulation techniques (Li et al., 2022; Liang et al., 2019) become requisite for such inverse problems. Prior arts are often designed for inverse problems of a handful of parameters. They either become prohibitive or do not converge when the inverse problem has a large number of unknown variables to be optimized. Deep neural networks offer a powerful modality for extracting knowledge from observations. They have also been applied for garment and cloth animation. For instance, Santesteban et al. (2019) show that deep nets can be exploited to predict garment deformation based on physics-based simulation results. PBNS (Bertiche et al., 2020) and SNUG (Santesteban et al., 2022) use unsupervised learning for the synthesis of garment deformation. Bertiche et al. (2022) further generalize this idea to learn dynamic garment movements. However, existing methods mostly focus on sheet-based models. A learning-based yarn-level model remains under-explored. This is likely due to the lack of high-quality training data and efficient simulation algorithms.
@@
An efficient forward simulation is often the key ingredient of the inverse problem. As the bottleneck in cloth simulation is often at solving the energy Hessian, a natural thought is to avoid a full linear solve in classic Newton’s method. Following this idea, Hecht et al. (2012) propose a lagged factorization scheme that reuses existing Cholesky factorization to save the computation. Multi-resolution (Capell et al., 2002; Lee et al., 2010) and multigrid (Tamstorf et al., 2015; Wang et al., 2018) solvers project fine-grid residual errors onto a coarser grid, on which linear or nonlinear iterations are more effective (Bolz et al., 2003; Tamstorf et al., 2015; Xian et al., 2019; Zhu et al., 2010). Liu et al. (2013) treat the implicit Euler integration as an energy minimization problem, in which spring constraints can be solved in parallel in the local step, while the global linear system remains constant on run-time. Projective Dynamics (PD) (Bouaziz et al., 2014) extend this concept to support a wider range of hyperelastic material models in quadratic form. Additionally, the computation of integration has been accelerated using methods such as the Chebyshev (Wang, 2015; Wang and Yang, 2016), Gauss-Seidel (Fratarcangeli et al., 2016), and various parallelization techniques on GPU (Wang and Yang, 2016; Fratarcangeli et al., 2016; Wu et al., 2020).
@@
Our work is also closely related to computational homogenization (Andreassen and Andreasen, 2014; Allaire and Brizzi, 2005; Grinspun et al., 2002) and numerical coarsening (Chen et al., 2017, 2018). This category of computing techniques aims to extract a low-rank representation of complex systems via global (coarse mesh) to local (fine mesh) optimization so that simulation at runtime efficiently captures the correct dynamics even on coarse grids (Kharevych et al., 2009; Nesme et al., 2009). It has been successfully applied for fabrication (Panetta et al., 2015; Chen et al., 2017) or for accelerated simulation (Torres et al., 2016). Chen et al. (2018) fit a material-aware shape function so that the coarse mesh replicates the behavior of a fine model. Constitutive model homogenization focuses more on high-level material responses (De Souza Neto et al., 2015; Blanco et al., 2016), where the underlying strain-stress relation is synthesized at REV (representative volume element). This method has also been applied to the inverse problem of micro-structure optimization or topology optimization (Eschenauer and Olhoff, 2001). For instance, Schumacher et al. (2015) build metamaterials with prescribed bulking behavior. Chen et al. (2015) propose a data-driven extension of this potential energy fitting idea to non-linear materials where now a coarse constitutive model is found through linear regression based on a set of deformation samples obtained from random forcing.
@@
Our method shares many similarities in algorithmic rationale and method design with those prior arts and specifically targeting realistic and efficient knitwear simulation. Our method is directly relevant and strongly inspired by recent contributions on HYLC (Sperl et al., 2020, 2022). HYLC shows a sheet-level homogenization paradigm that regresses a spline-based constitutive model. The homogenization substantially reduces the DOF, and the sheet-level simulation is much faster than YLS. The major difference between our method and HYLC is that our homogenization escalates the dimensionality of coarse mesh i.e., from a sheet- or rod-based discretization to a volumetric one. On the surface, such a strategy does not align with the original motivation of homogenization or coarsening, as it uses more DOF. Yet, we show that the increased dimensionality simplifies the material design when incorporated with a volume preservation constraint to homogenize bending and twisting effects of codimensional models. We are not the first to leverage volume preserving to model nonlinear bending deformation. Chen et al. (2023) use a volumetric prism element to simulate thick garments and have demonstrated the feasibility of this approach. Another major difference is that we learn a set of independent material parameters at each volume element, resulting in a high-dimensional space-time optimization problem. We name our method meta homogenization, hinting at these novel features of our formulation. The complexity of meta-homogenization is unlike most prior methods. To address this, we propose to leverage a domain-decomposed PD solver and adjoint Gauss-Newton to make the homogenization manageable. Meta homogenization is more expressive than HYLC–we can replicate the dynamics of complex and non-repetitious knits. It is also more efficient–the use of simpler material models with a fast solver offset the increase in homogenized DOF.
@@ 3. Method Overview
Figure 2. Meta-homogenization pipeline.  Given an input rod-based yarn model, our pipeline generates a volume mesh encapsulating the entire yarn structure. We lump the yarn-level mass to the mesh’s nodal points so that the inertia effect can be isolated. The yarn-to-mesh and mesh-to-yarn shape fitting facilitates us to define the loss function for homogenization. Per-element material parameters are obtained via an adjoint Gauss-Newton procedure, in a sample-by-sample manner. With a domain-decomposed PD solver, our method produces high-quality animation of knitwear garments that is visually similar to the full-scale YLS result. But we are two orders faster.
Our framework allows stable and efficient knitwear simulation that closely mimics the behavior of a full-scale YLS. To achieve this objective, our pipeline estimates a spatially varying homogenization scheme for the yarn fabric. Unlike prior methods that homogenize a triangular mesh approximating the fabric mid-surface (Sperl et al., 2020; Feng et al., 2024) and assign uniform material parameters across the mesh, our pipeline, as shown in Fig. 2, begins by constructing a volumetric mesh to encapsulate the given yarn-level model and computes the mass for each nodal point (Sec. 4). With the sequence of yarn-level simulated results, we proceed to compute the best-fitting material parameter for each volume element. This is a difficult task, because we do not have bounding volumetric mesh for each frame, the material parameters over mesh are of high dimension, and the optimization is sensitive to the initial condition. Therefore, we decompose the problem into two fitting steps. First, we formulate an optimization problem to find the best-fitting mesh shape matching the yarn-level deformation (Sec. 5). With the fitted shape, we employ an adjoint Gauss-Newton formulation that estimates a quasi-second-order descent direction of the high-dimension material parameter (Sec. 6). Adjoint Gauss-Newton synergizes with a harmonic initialization algorithm to progressively explore a good initial guess. To utilize the homogenized mesh generated from our fitting pipeline, we introduce a novel GPU-based forward simulator (Sec. 7). The global PD matrix is partitioned into domains corresponding to different knit patterns, and we leverage component mode synthesis (CMS) (Craig Jr and Hale, 1988) to build a subspace preconditioner for the global stage solve, followed by a full-space GPU-based Jacobi iteration.

### v3

# Volumetric Homogenization for Knitwear Simulation

@@ Abstract.
This paper presents volumetric homogenization, a spatially varying homogenization scheme for knitwear simulation. We are motivated by the observation that macro-scale fabric dynamics is strongly correlated with its underlying knitting patterns. Therefore, homogenization towards a single material is less effective when the knitting is complex and non-repetitive. Our method tackles this challenge by homogenizing the yarn-level material locally at volumetric elements. Assigning a virtual volume of a knitting structure enables us to model bending and twisting effects via a simple volume-preserving penalty and thus effectively alleviates the material nonlinearity. We employ an adjoint Gauss-Newton formulation(Zehnder et al., 2021) to battle the dimensionality challenge of such per-element material optimization. This intuitive material model makes the forward simulation GPU-friendly. To this end, our pipeline also equips a novel domain-decomposed subspace solver crafted for GPU projective dynamics, which makes our simulator hundreds of times faster than the yarn-level simulator. Experiments validate the capability and effectiveness of volumetric homogenization. Our method produces realistic animations of knitwear matching the quality of full-scale yarn-level simulations. It is also orders of magnitude faster than existing homogenization techniques in both the training and simulation stages.
submissionid: 379
@@
Figure 1. Yangge dance.  We propose a volumetric homogenization algorithm for knitwear simulation. Our pipeline takes full-scale yarn-level simulation observations as input and learns a homogenization out of the input poses. We name this procedure volumetric-homogenization because we do not homogenize yarn-level dynamics to a codimensional sheet/shell domain but to a volumetric enclosure. Doing so allows us to use an intuitive hyperelastic material model with a volume-preserving constraint to capture nonlinear behaviors like bending and twisting. volumetric homogenization varies spatially. Each volumetric element has its own material parameter, which substantially enhances the expressivity of our model. This is a challenging problem, we utilize the adjoint Gauss-Newton method(Zehnder et al., 2021) to enhance the convergence of the adjoint method. Additionally, we propose harmonic initialization, sample-based optimization, and a novel domain-decomposed projective dynamics solver, which significantly accelerates the entire optimization process. In the teaser figure, we show snapshots of an animated virtual character performing a Yangge dance, where we simulate a homogenized tetrahedron mesh with about $390$K elements that encapsulate a sweater. Cable patterns on the garment contribute to an irregular and sophisticated material distribution (as visualized in the sub-figure). In this example, the mesh-level simulation runs at $604$ ms per frame on average under $\Delta t=1/150$ sec. The yarn-level sweater model, on the other hand, consists of over $3$M DOFs. Our volumetric homogenization makes the simulation 100$\sim$ 100$\times$ 100 × faster than simulating the sweater at the yarn level.
[…]
@@ 1. Introduction
We propose an entirely different perspective for efficient simulation of knits that can even handle non-repeating knitting patterns. In contrast to traditional material homogenization methods or HYLC, we introduce a novel approach: volumemetric homogenization. We focus on studying the heterogeneity, i.e., the spatial variation of material parameters while keeping the constitutional relation simple. In particular, we incorporate a virtual volume representing the yarn material and exploit spatially varying volume-preserving penalties to capture desired knit characteristics, such as stretching, shearing, and bending, without over-complicating the material model.
We discretize the fabric at the yarn level and embed it within a volumetric mesh. This mesh is of high resolution to capture the local structural characteristics of yarn loops. We learn a macro-scale material model over this volumetric space, aligning it with dynamic YLS results. In that respect, our method can also be understood as applying local homogenization at each mesh element. Our rationale resembles curve fitting, where using multiple low-degree polynomials (e.g., splines) is often preferred over employing a single high-degree polynomial. Likewise, a composite material model, despite its simplicity, offers a more versatile design space compared to a homogenized material, since each volumetric element possesses independent material freedoms. We employ the adjoint Gauss-Newton method(Zehnder et al., 2021) to regress the yarn material via a high-dimension space-time optimization problem. The simplicity of the material model makes the volumetric homogenization well-suited for efficient forward GPU solvers such as projective dynamics (PD) (Bouaziz et al., 2014). To this end, we devise a domain-decomposed multi-level solver for the global stage of PD specially crafted for our runtime pipeline.
@@ 2. Related work
Modern fabrics and garments are fabricated through sophisticated workmanship and exhibit intricate material response. Nevertheless, modeling cloth with a sheet-like shell remains prevalent (Choi and Ko, 2002; Grinspun et al., 2003) due to its simplicity and intuitiveness. Mass-spring and/or triangle meshes offer a natural discretization of sheet-based cloth, and its equation of motion can be derived via elastic energies that measure the cloth deformation (Terzopoulos et al., 1987). Implicit integration has become a standard component since the seminal work of Baraff and Witkin (1998). Unlike a rubber membrane, cloth fabrics are less stretchable but tend to be bent more easily. This feature inspires various techniques to enforce the inextensibility (Provot, 1995; Goldenthal et al., 2007; English and Bridson, 2008; Wang et al., 2010) with numerical stability. Cloth bending, on the other hand, can be parameterized by the dihedral angle or bending modes (Bridson et al., 2003). It is also possible to exploit isometric mesh deformations to use the so-called quadratic bending (Bergou et al., 2006; Garg et al., 2007), which possesses a constant Hessian. Kim (2020) reveals the connection between model of Baraff and Witkin (1998) and the finite element method (FEM) (Bathe, 2006). To capture detailed wrinkles and folds on the cloth, adaptive remeshing (Narain et al., 2013, 2012) or mixed discretization (Guo et al., 2018; Weidner et al., 2018) are proven to be effective. Another important aspect of sheet-level cloth animation is to handle collisions, especially self-collisions (Volino and Thalmann, 2000; Bridson et al., 2002; Baraff et al., 2003). Penalty-based methods (Bouaziz et al., 2014; Wu et al., 2020) are straightforward to implement but often with stability issues. Provot (1997); Harmon et al. (2008) introduce approaches that group multiple collisions into “impact zones” and treat them as rigid bodies, allowing for some sliding motion. Constraint-based collision handling methods represent contact as constraints based on exact Coulomb friction (Otaduy et al., 2009; Li et al., 2018).
@@
Data-driven or learning-based cloth simulation is considered an effective approach to enhance the realism of continuum shell models. Wang et al. (2011) estimate planar and bending stiffness through separate tests, while automatic devices have been developed for acquiring fabric stiffness (Miguel et al., 2012) and friction parameters (Miguel et al., 2013). Similar tensile tests have also been utilized by Clyde et al. (2017) to estimate the planar stiffness of woven fabrics. Feng et al. (2022) use a simulation-in-the-loop framework to estimate the fabric bending stiffness. ClothCap aims to reconstruct multiple garments from the 3D scans (Pons-Moll et al., 2017). Differentiable simulation techniques (Li et al., 2022; Liang et al., 2019) become requisite for such inverse problems. Prior arts are often designed for inverse problems of a handful of parameters. They either become prohibitive or do not converge when the inverse problem has a large number of unknown variables to be optimized. Deep neural networks offer a powerful modality for extracting knowledge from observations. They have also been applied for garment and cloth animation. For instance, Santesteban et al. (2019) show that deep nets can be exploited to predict garment deformation based on physics-based simulation results. PBNS (Bertiche et al., 2021) and SNUG (Santesteban et al., 2022) use unsupervised learning for the synthesis of garment deformation. Bertiche et al. (2022) further generalizes this idea to learn dynamic garment movements. However, existing methods mostly focus on sheet-based models. A learning-based yarn-level model remains under-explored. This is likely due to the lack of high-quality training data and efficient simulation algorithms.
@@
An efficient forward simulation is often the key ingredient of the inverse problem. As the bottleneck in cloth simulation is often at solving the energy Hessian, a natural thought is to avoid a full linear solve in classic Newton’s method. Following this idea, Hecht et al. (2012) propose a lagged factorization scheme that reuses existing Cholesky factorization to save the computation. Multi-resolution (Capell et al., 2002; Lee et al., 2010) and multigrid (Tamstorf et al., 2015; Wang et al., 2018) solvers project fine-grid residual errors onto a coarser grid, on which linear or nonlinear iterations are more effective (Bolz et al., 2003; Tamstorf et al., 2015; Xian et al., 2019; Zhu et al., 2010). Liu et al. (2013) treat the implicit Euler integration as an energy minimization problem, in which spring constraints can be solved in parallel in the local step, while the global linear system remains constant on run-time. Projective Dynamics (PD) (Bouaziz et al., 2014) extend this concept to support a wider range of hyperelastic material models in quadratic form. Additionally, the computation of integration has been accelerated using methods such as the Chebyshev (Wang, 2015; Wang and Yang, 2016), Gauss-Seidel (Fratarcangeli et al., 2016), and various parallelization techniques on GPU (Wang and Yang, 2016; Fratarcangeli et al., 2016; Wu et al., 2020). Fully leveraging the obtained gradient to find a better search direction to improve convergence is also essential for achieving better performance. A common strategy involves estimating the second-order information of the system, e.g. L-BFGS(Du et al., 2021; Li et al., 2022) or Anderson acceleration(Peng et al., 2018). Wang (2018) and Zimmermann et al. (2019) propose using the Gauss-Newton solver, which exhibits promising convergence. However, assembling and factorizing the dense Gauss-Newton solver becomes a new bottleneck. Similar to how the adjoint method circumvents dense matrices, Zehnder et al. (2021) introduces two additional adjoint variables that transform the Gauss-Newton solver to be a sparse linear system, greatly accelerating the performance.
Figure 2. Volumetric homogenization pipeline.  Given an input rod-based yarn model, our pipeline generates a volume mesh encapsulating the entire yarn structure. We lump the yarn-level mass to the mesh’s nodal points so that the inertia effect(Eq. 14) can be isolated. The yarn-to-mesh and mesh-to-yarn shape fitting facilitates us to define the loss function for homogenization. Per-element material parameters are obtained via an adjoint Gauss-Newton procedure, in a sample-by-sample manner. With a domain-decomposed PD solver, our method not only produces high-quality animation of knitwear garments that is visually similar to the full-scale YLS result but also achieves these results two orders of magnitude faster.
[…]
@@
Our work is also closely related to computational homogenization (Andreassen and Andreasen, 2014; Allaire and Brizzi, 2005; Grinspun et al., 2002) and numerical coarsening (Chen et al., 2017, 2018). This category of computing techniques aims to extract a low-rank representation of complex systems via global (coarse mesh) to local (fine mesh) optimization so that simulation at runtime efficiently captures the correct dynamics even on coarse grids (Kharevych et al., 2009; Nesme et al., 2009). It has been successfully applied for fabrication (Panetta et al., 2015; Chen et al., 2017) or for accelerated simulation (Torres et al., 2016). Chen et al. (2018) fit a material-aware shape function so that the coarse mesh replicates the behavior of a fine model. Constitutive model homogenization focuses more on macroscopic material responses (De Souza Neto et al., 2015; Blanco et al., 2016), where the underlying strain-stress relation is synthesized at RVE (representative volume element). This method has also been applied to the inverse problem of micro-structure optimization or topology optimization (Eschenauer and Olhoff, 2001). For instance, Schumacher et al. (2015) build deformable objects with target stiffness using cells of prescribed bulking behavior. Chen et al. (2015) propose a data-driven extension of this potential energy fitting idea to non-linear materials where now a coarse constitutive model is found through linear regression based on a set of deformation samples obtained from random forcing.
@@
Our method shares many similarities in algorithmic rationale and method design with those prior arts and specifically targeting realistic and efficient knitwear simulation. Our method is directly relevant and strongly inspired by recent contributions on HYLC (Sperl et al., 2020, 2022). HYLC shows a sheet-level homogenization paradigm that regresses a spline-based constitutive model. The homogenization substantially reduces the DOF, and the sheet-level simulation is much faster than YLS. The major difference between our method and HYLC is that our homogenization escalates the dimensionality of coarse mesh i.e., from a sheet- or rod-based discretization to a volumetric one. On the surface, such a strategy does not align with the original motivation of homogenization or coarsening, as it uses more DOF. Yet, we show that the increased dimensionality simplifies the material design when incorporated with a volume preservation constraint to homogenize bending and twisting effects of codimensional models. We are not the first to leverage volume preserving to model nonlinear bending deformation. Chen et al. (2023) use a volumetric prism element to simulate thick garments and have demonstrated the feasibility of this approach. Another major difference is that we learn a set of independent material parameters at each volume element, resulting in a high-dimensional space-time optimization problem. We name our method volumetric homogenization, hinting at these novel features of our formulation. The complexity of volumetric homogenization is unlike most prior methods. To address this, we propose to leverage a domain-decomposed PD solver and employ adjoint Gauss-Newton(Zehnder et al., 2021) to make the homogenization manageable. Volumetric homogenization is more expressive than HYLC–we can replicate the dynamics of complex and non-repetitious knits. It is also more efficient–the use of simpler material models with a fast solver offsets the increase in homogenized DOF.
@@ 3. Method Overview
Our framework allows stable and efficient knitwear simulation that closely mimics the behavior of a full-scale YLS. To achieve this objective, our pipeline estimates a spatially varying homogenization scheme for the yarn fabric. Unlike prior methods that homogenize a triangular mesh approximating the fabric mid-surface (Sperl et al., 2020; Feng et al., 2024) and assign uniform material parameters across the mesh, our pipeline, as shown in Fig. 2, begins by constructing a volumetric mesh to encapsulate the given yarn-level model and computes the mass for each nodal point (Sec. 4). With the sequence of yarn-level simulated results, we proceed to compute the best-fitting material parameter for each volume element. This is a difficult task, because we do not have bounding volumetric mesh for each frame, the material parameters over mesh are of high dimension, and the optimization is sensitive to the initial condition. Therefore, we decompose the problem into two fitting steps. First, we formulate an optimization problem to find the best-fitting mesh shape matching the yarn-level deformation (Sec. 5). With the fitted shape, we employ an adjoint Gauss-Newton formulation(Zehnder et al., 2021) that estimates a quasi-second-order descent direction of the high-dimension material parameter (Sec. 6). Adjoint Gauss-Newton synergizes with a harmonic initialization algorithm to progressively explore a good initial guess. To utilize the homogenized mesh generated from our fitting pipeline, we introduce a novel GPU-based forward simulator (Sec. 7). The global PD matrix is partitioned into domains corresponding to different knit patterns, and we leverage component mode synthesis (CMS) (Craig Jr and Hale, 1988) to build a subspace preconditioner for the global stage solve, followed by a full-space GPU-based Jacobi iteration.